# Liouvillian Gap in Dissipative Haar-Doped Clifford Circuits

### v1

@@ Abstract
Quantum chaos is commonly assessed through probe-dependent signatures such as spectral statistics, OTOCs, and entanglement growth, which need not coincide. Recently, a dissipative diagnostic of chaos has been proposed, in which an infinitesimal coupling to a bath yields a finite Liouvillian gap in chaotic systems, marking the onset of intrinsic relaxation. This raises a conceptual question: what is the minimal departure from Clifford dynamics needed for this intrinsically relaxing behavior to emerge? In this work, we investigate the dynamics under the Floquet two-qubit Clifford circuit interleaved with a finite density of Haar-random single-site gates, followed by a depolarizing channel with strength $\gamma$. For Floquet Clifford circuits built from an iSWAP-class two-qubit gate, our analysis identifies two distinct regimes for the Liouvillian gap in the thermodynamic limit, exemplified by the undoped and fully doped extreme cases. In both regimes, the dissipative diagnostic signals chaotic behavior, differing only in how the gap scales with system size. In the undoped circuit, the gap scales as $\Delta\sim\gamma N$, whereas in the fully doped circuit it remains finite as $N\to\infty$. We find that the doping density $p_{h}$ governs the crossover: as $p_{h}\to 0$, any spatial structure remains undoped-like, whereas for finite $p_{h}$ certain structures can enter a finite-gap regime. These results are analytically established in the strongly dissipative regime $\gamma\gg 1$ by deriving lower bounds on the gap as a function of $p_{h}$ and explicit finite-gap constructions, and their extension toward $\gamma\to 0$ is supported by numerics. Importantly, our analytic treatment depends only on the spatial doping structure, so the same gap scaling persists even when the Haar rotations are independently resampled each Floquet period.
@@
Figure 1: Summary of analytic results. Dissipative Haar-doped Floquet circuit and the weight-truncation framework for bounding the Liouvillian gap. (a) One Floquet period consists of a two-layer Clifford brickwork built from a fixed iSWAP-class gate on odd and even bonds (blue), with single-qubit Haar rotations applied immediately after each layer on a chosen set of $n_{h}$ qubit lines (orange), followed by an onsite depolarizing layer on all qubits (yellow). In (b)-(d), each long rectangle represents the qubit lines, with blue and orange segments indicating undoped and doped sites, respectively. (b) Diagram of Liouvillian gap $\Delta$ versus dissipation strength $\gamma$ for representative doping patterns at the thermodynamic limit. Solid lines show exact results for fully doped (orange); dashed lines show analytic upper bounds for dense doping (light purple) and block-staggered doping with spacing $k$ between doped (green). Unlike the doped patterns, whose singularities remain independent of system size, the undoped circuit (blue) exhibits a singularity that diverges as the system size NN increases. (c,d) Lower- and upper-bound constructions. Gray capacitors denote Pauli-weight truncation projectors inserted between Floquet steps. (c) Representative truncation of an initially local $Y$ operator (green) once its weight exceeds $w_{\mathrm{t}}=6$ and vanishes after five steps. (d) Under truncation, return cycles can only be supported by Pauli strings confined to a small consecutive block, illustrated by a three-site string $YXZ$ that translates by two sites per step and returns after $N/2$ steps.
@@ I Introduction
In classical mechanics, chaos is defined at the level of trajectories based on the sensitivity to initial conditions. In quantum many-body systems, no single definition plays the same role; quantum chaos is typically inferred from a collection of signatures, including spectral statistics [haake1991quantum, bohigas1984characterization, guhr1998random, d2016quantum, deutsch2018eigenstate, eisert2015quantum, leblond2021universality], out-of-time-order correlation (OTOC) based measures of operator growth/scrambling [roberts2017chaos, harrow2021separation, hosur2016chaos, nahum2018operator], or entanglement growth diagnostics [prosen2007efficiency, nahum2017quantum]. However, these signatures can disagree [dowling2023scrambling, von2018operator, turner2018weak, pizzi2025genuine], so no single probe is universally definitive. Recent work proposes an intrinsic relaxation rate as follows [mori2024liouvillian]: given a Lindbladian dynamics with Hamiltonian $H$ and dissipation parametrized by $\gamma$, take the thermodynamic limit $N\rightarrow\infty$ first, and then let $\gamma\rightarrow 0^{+}$, and ask whether the leading relaxation rate, namely Liouvillian gap, remains finite. In settings without conservation laws or other hydrodynamic slow modes, a finite intrinsic relaxation rate is expected due to rapid operator spreading that makes dissipation act on an extensive operator support. Such an operator spreading timescale can also be accessed in the frequency domain via dissipative spectral form factors and related Ruelle-Pollicott-type resonance structures [liao2022emergence, yoshimura2024robustness, yoshimura2025theory, zhang2024thermalization, jacoby2025spectral]. In contrast, in integrable settings or in systems with (quasi-)conservation, the weak-dissipation singularity can be suppressed or take a qualitatively different form [garcia2023keldysh, vznidarivc2024momentum, duh2025ruelle].
@@
Motivated by this tension, we start from a Clifford backbone and introduce non-Clifford doping as a controlled departure. Doping injects non-stabilizer resources, allowing one to interpolate between a stabilizer circuit ensemble and a Haar-random one. In many works, this interpolation is tracked through spectral and eigenstate diagnostics, which cross over toward random-matrix universality; importantly, the doping scale required is probe dependent. For instance, spectral statistics can cross over at ${\cal O}(1)$ doping [zhou2020single], whereas higher-point OTOC-based diagnostics may require resources scaling extensively with system size [leone2021quantum]. Instead, in this work, we apply the intrinsic gap diagnostic to understand dissipative Haar-doped Floquet Clifford circuits, illustrated in Fig. 1(a) and detailed in Sec. II. We find that even the undoped circuit, for a suitable iSWAP-class Clifford gate, exhibits a pronounced weak-dissipation singularity of the Liouvillian gap in the thermodynamic limit. Moreover, when the doping density is nonvanishing, i.e., the number of doped qubit lines scales proportionally with system size $N$, the system-size scaling of this singular response crosses over: an $N$-independent scaling emerges at finite doping density, replacing the $N$-growing behavior of the undoped circuit. Importantly, this crossover is insensitive to whether the Haar rotations are quenched or independently redrawn each Floquet period. We elaborate on these findings below by summarizing our main analytic results and supporting numerics.
I.1 Summary of the Result
We begin with the undoped Clifford limit, which serves as a baseline for our diagnostic. In Sec. III, we find that the undoped iSWAP-class Floquet circuit exhibits a deterministically emerging weak dissipation singularity of the Liouvillian gap. The other locally Clifford-equivalent two-qubit gate classes do not show the same singular response. Moreover, unlike commonly studied chaotic Floquet circuits [yoshimura2024robustness], the gap in this undoped iSWAP-class circuit grows proportionally with system size $N$, and therefore appears to diverge in the thermodynamic limit. In this sense, the iSWAP-class circuit already provides a strongly chaos-motivated reference point within a purely Clifford, classically simulable setting. For analytic concreteness, we fix a representative iSWAP-class Clifford backbone throughout (see Sec. II).
When we consider nonzero doping density, we can no longer rely on purely Clifford dynamics to analytically compute the gap at any finite $\gamma$. Therefore, in Sec. IV, we turn to numerical simulation to characterize the Liouvillian gap in the weak dissipation regime. We argue that the observed linear growth of $\langle\Delta\rangle$ with $N$ indicates a weak dissipation singularity, which can then be interpreted as evidence of chaotic dynamics (see Fig. 2(a,b) for representative examples). This observation aligns with our expectation that chaotic behavior should not be eliminated by introducing randomness to the dynamics via Haar-doping.
To understand when this doping-induced modification leads to a crossover in system-size scaling, we turn to the strong-dissipation regime, which provides an analytically tractable limit. Building on the effective Pauli-dynamics description introduced in Sec. II.4 and developed further in Sec. V, we obtain the following lower bound on the Liouvillian gap:
@@ I.1 Summary of the Result
A precise statement and proof are given in Theorem 1. The proof is based on an effective Pauli-dynamics picture in the strong dissipation regime: because depolarizing noise exponentially suppresses high Pauli weight, we can track whether operator trajectories can stay within a low-weight sector over many Floquet steps. If every trajectory starting in this sector is eventually driven above the cutoff, then the truncated dynamics supports no long-time decay modes, forcing any slow relaxation of the full channel to proceed through excursions to higher weight and yielding a lower bound on the Liouvillian gap. As illustrated in Fig. 1(c), under Floquet steps a low-weight Pauli string typically spreads through undoped regions, so its Pauli weight is driven above the cutoff and it is removed in the truncated description.
This lower bound already captures the essential effect of sparse doping on the system-size scaling of the Liouvillian gap. If the doping density $n_{h}/N$ vanishes as $N\to\infty$, it forces the gap to grow with system size for any such doping structure. When instead $n_{h}/N$ remains finite, the bound does not enforce such divergence. In Sec. VI, we demonstrate how representative doping structures can yield a finite Liouvillian gap in the thermodynamic limit.
As an opposite extreme to the undoped circuit, the fully doped pattern provides a natural reference point in which local Haar mixing is maximized. In this case the gap is obtained exactly as $\Delta=\gamma+2$. See Theorem 3. We further construct doped patterns with nonvanishing doping density for which the gap can remain finite as $N\to\infty$, as summarized in Fig. 1(b). For dense doping, where every undoped site is separated by at least one doped site, we derive the explicit upper bound $\Delta\leq 3\gamma+3$; see Theorem 4. For block-staggered doping, doped sites are separated by $k$ consecutive undoped sites, providing a structure with nonvanishing doping density that can nevertheless approach a sparse limit as $k$ increases. For this structure we obtain an upper bound $\Delta\leq A\gamma+B$ with $A$ and $B$ growing at most linearly in $k$; see Sec. VI.3. In particular, for any fixed $k$ independent of system size, the gap remains finite at the thermodynamic limit.
The basic idea behind our upper bounds is to identify return cycles that survive within the effective Pauli dynamics, since such cycles generate long-time decay modes in the truncated description. As illustrated in Fig. 1(d), the surviving contributions come from translation-like return paths confined to a small block of sites, which in turn yields an upper bound on the Liouvillian gap (see Sec. VI.2). By contrast, Pauli strings that are not confined to such a block typically spread ballistically, with the two endpoints expanding in opposite directions. As this growing support repeatedly traverses undoped regions, Clifford propagation drives the Pauli weight upward until it exceeds the cutoff and is removed by truncation.
@@
Taken together, our results identify two distinct regimes for the Liouvillian gap in the thermodynamic limit, exemplified by the undoped and fully doped extreme cases. In the undoped circuit the singular response grows with system size, whereas in the fully doped circuit it remains finite as $N\to\infty$. When $n_{h}/N\to 0$, any doping structure remains in the undoped-like regime with system-size growth, whereas for nonvanishing $n_{h}/N$ a crossover becomes possible and certain structures enter the fully-doped-like regime with a finite gap in the thermodynamic limit. We close by emphasizing a notable consequence of the way we derive our upper and lower bounds Because the construction of our upper and lower bounds depends only on the spatial doping pattern, it does not rely on reusing the same single-qubit Haar rotations from one Floquet period to the next. Accordingly, the same picture applies both for quenched Haar disorder and for Haar rotations resampled independently each Floquet period.
@@ II.2 Floquet Channel and Liouvillian Gap
Let $e^{\lambda_{a}}$ denote the eigenvalues of $\Phi_{F}$, with $\lambda_{0}=0$ corresponding to the unique periodic steady state. Uniqueness follows from the depolarizing stage, which mixes any initial state toward the maximally mixed state and hence renders the one-period channel primitive. We refer to $\{\lambda_{a}\}$ as the Liouvillian eigenvalues in the Floquet setting. The Liouvillian gap is defined as
@@ III.2.2 Eigenspectrum dependence on localization structure
Therefore, in the absence of doping, the Liouvillian gap in Eq. (27) grows linearly in $N$ for all $\gamma$, so already the thermodynamic limit $N\rightarrow\infty$ and weak dissipation limit $\gamma\rightarrow 0$ do not commute. From the viewpoint of dissipative diagnostics of quantum chaos, the underlying dynamics of this dissipative Clifford circuit can be seen as maximally chaotic (see Sec. VII) according to the dissipative diagnostic. In some sense, this is surprising as Clifford circuits are deemed to be not complex due to the fact that they are efficiently simulated in classical computers, which is at odds with our explicit derivation that the system is maximally chaotic in a Liouvillian gap diagnostic.
@@ IV Weak Dissipation Singularity
We argue that the numerical results presented in Fig. 2, in conjunction with Result 1, imply a singularity in the Liouvillian gap in the thermodynamic limit for any nonzero $p_{h}$. For such doping patterns, our expectation is that as the maximum system size $N$ increases, the threshold $\gamma^{*}$ below which the gap scales linearly in $N$ will decrease. Then, at any nonzero $\gamma$, the gap saturates to a finite value as $N$ increases, so $\Delta_{\gamma\rightarrow+0}$ will retain a finite value in the thermodynamic limit (but $\Delta_{\gamma=0}=0$, because $\gamma=0$ corresponds to the unitary case in which the eigenvalues of the Floquet channel have modulus 1).
@@ VII Discussion
First, our results indicate, according to the dissipative diagnostic, that the circuit without doping is maximally chaotic in a concrete sense: for any fixed, arbitrarily small $\gamma>0$, the Liouvillian gap increases with system size, so the approach to the steady state becomes faster (rather than slower) as $N$ grows. The mechanism is simple. The unitary dynamics rapidly spreads typical local Pauli operators to extensive weight $w(t)={\cal O}(N)$ on an ${\cal O}(N)$ timescale [Spreading2018]. Once operators are extensive, a local noise channel of strength $\gamma$ damps them at a rate $\sim\gamma w$, i.e. $\sim\gamma N$. In the absence of conserved densities or other slow channels, there is no hydrodynamic mechanism that can produce small decay rates, and the Liouvillian spectrum develops a parametrically large gap.
Second, adding Haar random gates, contrary to our initial intuition based on simulation complexity, makes the circuit less chaotic in the sense that the intrinsic relaxation does not scale with system size upon doping. At the level of Pauli dynamics, local randomness induces strong Pauli-weight mixing that regenerates low-weight components at each Floquet step, a process that is precluded in the undoped circuit (which is fine-tuned). These low-weight components dominate the slowest nontrivial eigenmodes and set the intrinsic relaxation rate, thereby rendering the gap finite in the thermodynamic limit. Our theorem precisely shows that this crossover from unbounded growth of the gap with system size to a finite Liouvillian gap sets in once the doping density is nonvanishing, i.e., when the number of doped sites scales linearly with system size. Notably, the scale of doping density at which this crossover sets in coincides with the regime in which scrambling diagnostics (e.g., OTOCs) become Haar-like [leone2021quantum, magni2025quantum] and higher $t$-moments approach those of Haar-random unitaries [haferkamp2022random, harrow2023approximate]. In this sense, the appearance of a finite Liouvillian gap tracks the same extensive-doping regime that drives Haar-like behavior in standard chaos diagnostics.
@@
Taken together, our finding that a finite gap sets in once the number of doped sites scales linearly with system size and prior results on Haar-like scrambling point to thresholds of the same order for relaxation- and scrambling-based diagnostics. Interpreting this comparison, however, requires some care. Most existing results are established in circuit ensembles with a different microscopic architecture, typically at greater depth or in different geometries than the shallow one-dimensional setting analyzed here. Bringing these diagnostics into a single controlled framework therefore calls for extending the present model beyond one dimension and beyond the shallow-depth setting. An important direction for future work is to connect the finite-gap scale to approximate unitary $t$-design formation, thereby clarifying how these two diagnostics are related. Ultimately, establishing such links between complementary diagnostics may help develop a more unified picture of quantum chaos beyond diagnostic-specific characterizations.

### v2

@@ Abstract
Quantum chaos is commonly assessed through probe-dependent signatures that need not coincide. Recently, a dissipative signature was proposed for chaotic Floquet systems, where infinitesimal bulk dissipation induces a non-zero constant intrinsic relaxation rate quantified by the Liouvillian gap. This raises a question: what minimal departure from Clifford dynamics is required to generate such intrinsic relaxation? To address this, we study a Floquet two-qubit Clifford circuit doped with Haar-random single-qubit gates and subject to local dissipation of strength $\gamma$. We find a structure-dependent crossover. The undoped iSWAP-class circuit exhibits a weak-dissipation singularity, with a gap that grows with $N$ for any $\gamma>0$. Haar doping preserves this undoped-like growth for any subextensive doping pattern. At finite doping density, there exist patterns that yield an $\mathcal{O}(1)$ gap for any fixed $\gamma$ as $N\to\infty$, yet remain singular as $\gamma\to 0^{+}$. Because our bounds depend only on the spatial doping pattern, they remain valid even when the Haar rotations are independently redrawn each Floquet period. Overall, our findings provide a circuit-level perspective on intrinsic relaxation—and thus irreversibility—in open many-body systems.
@@
Figure 1: Summary of analytic results. Dissipative Haar-doped Floquet circuit and the weight-truncation framework for bounding the Liouvillian gap. (a) One Floquet period consists of a two-layer Clifford brickwork built from a fixed iSWAP-class gate on odd and even bonds (blue), with single-qubit Haar rotations applied immediately after each layer on a chosen set of $n_{h}$ qubit lines (orange), followed by an onsite depolarizing layer on all qubits (yellow). In (b)-(d), each long rectangle represents the qubit lines, with blue and orange segments indicating undoped and doped sites, respectively. (b) Diagram of Liouvillian gap $\Delta$ versus dissipation strength $\gamma$ for representative doping patterns at the thermodynamic limit. Solid lines show exact results for fully doped (orange); dashed lines show analytic upper bounds for dense doping (light purple) and block-staggered doping with spacing $k$ between doped (green). Unlike the doped patterns, whose singularities remain independent of system size, the undoped circuit (blue) exhibits a singularity that diverges as the system size $N$ increases. (c,d) Lower- and upper-bound constructions. Gray capacitors denote Pauli-weight truncation projectors inserted between Floquet steps. (c) Representative truncation of an initially local $Y$ operator (green) once its weight exceeds $w_{\mathrm{t}}=6$ and vanishes after five steps. (d) Under truncation, return cycles can only be supported by Pauli strings confined to a small consecutive block, illustrated by a three-site string $YXZ$ that translates by two sites per step and returns after $N/2$ steps.
@@ I Introduction
In classical mechanics, chaos is defined at the level of trajectories based on the sensitivity to initial conditions. In quantum many-body systems, no single definition plays the same role; quantum chaos is typically inferred from a collection of signatures, including spectral statistics [haake1991quantum, bohigas1984characterization, guhr1998random, d2016quantum, deutsch2018eigenstate, eisert2015quantum, leblond2021universality], out-of-time-order correlation (OTOC) based measures of operator growth/scrambling [roberts2017chaos, harrow2021separation, hosur2016chaos, nahum2018operator], or entanglement growth diagnostics [prosen2007efficiency, nahum2017quantum]. However, these signatures can disagree [dowling2023scrambling, von2018operator, turner2018weak, pizzi2025genuine], so no single probe is universally definitive.
The emergence of irreversibility from pure Hamiltonian dynamics is often discussed using concepts from quantum chaos, in particular the eigenstate thermalization hypothesis (ETH) [deutsch1991quantum, srednicki1994chaos, rigol2008thermalization, mori2018thermalization]. Beyond this unitary perspective, recent work proposes an intrinsic relaxation signature for chaotic Floquet systems [mori2024liouvillian]. Specifically, for a Lindbladian dynamics with dissipation strength $\gamma$, one takes system size $N\to\infty$ first and then $\gamma\to 0^{+}$, and asks whether the Liouvillian gap remains finite. This weak-dissipation perspective has since been developed and extended in a variety of settings [yoshimura2024robustness, yoshimura2025theory, zhang2024thermalization, jacoby2025spectral]. Separately, spectral form factor diagnostics have been used to probe many-body chaos through controlled nonunitary deformations [liao2022emergence].
This signature is tied to operator spreading: without conservation laws or other hydrodynamic slow modes, rapid spreading makes local dissipation effectively act on an extensive support, yielding a finite intrinsic gap. By contrast, integrability or (quasi-)conservation can suppress this weak-dissipation response or alter it qualitatively [garcia2023keldysh, vznidarivc2024momentum, duh2025ruelle].
@@
Motivated by this tension, we start from a Clifford backbone and introduce non-Clifford doping as a controlled departure. Doping injects non-stabilizer resources, allowing one to interpolate between a stabilizer circuit ensemble and a Haar-random one. In many works, this interpolation is tracked through spectral and eigenstate diagnostics, which cross over toward random-matrix universality; importantly, the doping scale required is probe dependent. For instance, entanglement spectral statistics can cross over at ${\cal O}(1)$ doping [zhou2020single], whereas higher-point OTOC-based diagnostics may require resources scaling extensively with system size [leone2021quantum]. Instead, in this work, we apply the intrinsic gap diagnostic to understand dissipative Haar-doped Floquet Clifford circuits, illustrated in Fig. 1(a) and detailed in Sec. II. We summarize our main results below.
I.1 Summary of Results
We begin with the undoped Clifford limit as a baseline. In Sec. III, we find that the undoped iSWAP-class Floquet circuit has a Liouvillian gap that scales as $\Delta\propto N$ for any dissipation strength $\gamma$, whereas other locally Clifford-equivalent two-qubit gate classes do not. For concreteness, we fix a representative iSWAP-class Clifford gate throughout. Once we introduce Haar doping, the dynamics is no longer purely Clifford, so the Liouvillian gap is not exactly tractable at finite $\gamma$ by stabilizer methods alone. In Sec. IV, we use numerics to probe the weak-dissipation regime and find the Haar-averaged gap scales as $\langle\Delta\rangle\sim N$ (see Fig. 2(a)–(c)). This scaling indicates that the weak-dissipation response persists under Haar doping.
To understand the crossover in system-size scaling away from the $\gamma\to 0^{+}$ limit, we turn to the analytically tractable strong-dissipation regime. In this regime, the effective Pauli dynamics admits a natural description in terms of a weight truncation with a cutoff set by the dissipation strength. In Sec. V, we first obtain the following lower bound on the Liouvillian gap:
@@ I.1 Summary of Results
This bound can be viewed as a minimal-cutoff condition: if all trajectories initiated below the cutoff are eventually driven above it by Pauli-weight growth through undoped regions (see Fig. 1(c)), then the Liouvillian gap cannot be captured within the truncated sector alone. A precise statement and proof are given in Theorem 1. This lower bound captures the qualitative role of doping density. If $n_{h}/N\to 0$ as $N\to\infty$, it forces the gap to grow with system size, whereas at finite doping density it does not. In Sec. VI, we show that at nonvanishing doping density the Liouvillian gap can remain finite for suitable doping patterns as depicted in Fig. 1(b).
As an opposite extreme to the undoped circuit, the fully doped pattern provides a natural reference point with maximal local Haar mixing, for which the gap is obtained as $\Delta=\gamma+2$ (Theorem 3). For dense doping (no adjacent undoped sites), we derive an explicit upper bound $\Delta\leq 3\gamma+3$ (Theorem 4), while for block-staggered doping with $k$ consecutive undoped sites between doped sites we obtain $\Delta\leq A\gamma+B$ with $A,B=O(k)$, implying an $N$-independent gap for any fixed $k$ (see Sec. VI.3). These upper bounds follow from how trajectories behave under the truncated effective Pauli dynamics: ballistically spreading strings inevitably traverse an undoped region, where deterministic Clifford growth pushes the Pauli weight above the cutoff and the path is removed by truncation. By contrast, the surviving contributions are return cycles confined to a small contiguous block that evolve essentially by directed translation and close after $N/2$ steps (see Fig. 1(d)).
@@
Taken together, our results identify two thermodynamic regimes for the Liouvillian gap: undoped-like system-size growth versus a fully-doped-like $O(1)$ gap. If $n_{h}/N\to 0$, the lower bound forces growth for any pattern, whereas at nonvanishing doping density a crossover becomes possible and suitable structures can sustain a finite gap as $N\to\infty$. Because our bounds depend only on the spatial doping pattern, the same picture applies both to quenched Haar disorder and to Haar rotations resampled independently each Floquet period.
@@ II.2 Floquet Channel and Liouvillian Gap
Let $e^{\lambda_{a}\tau}$ denote the eigenvalues of $\Phi_{F}$, with $\lambda_{0}=0$ corresponding to the unique periodic steady state. Uniqueness follows from the depolarizing stage, which mixes any initial state toward the maximally mixed state and hence renders the one-period channel primitive. We refer to $\{\lambda_{a}\}$ as the Liouvillian eigenvalues in the Floquet setting. The Liouvillian gap is defined as
@@ III.2.2 Eigenspectrum dependence on localization structure
Therefore, in the absence of doping, the Liouvillian gap in Eq. (27) grows linearly in $N$ for any $\gamma$, so already the thermodynamic limit $N\rightarrow\infty$ and weak dissipation limit $\gamma\rightarrow 0^{+}$ do not commute. Under the gap diagnostic, the undoped Clifford circuit exhibits the strongest intrinsic relaxation in our setting. In some sense, this is surprising because Clifford circuits are classically efficiently simulable. The resolution is that the Liouvillian gap probes dissipative relaxation rather than unitary circuit complexity.
@@ IV Weak Dissipation Singularity
We argue that the numerical results presented in Fig. 2, in conjunction with Result 1, imply a singularity in the Liouvillian gap in the thermodynamic limit for any nonzero $p_{h}$. For such doping patterns, our expectation is that as the maximum system size $N$ increases, the threshold $\gamma^{*}$ below which the gap scales linearly in $N$ will decrease. Then, at any nonzero $\gamma$, the gap saturates to a finite value as $N$ increases, so $\Delta_{\gamma\rightarrow 0^{+}}$ will retain a finite value in the thermodynamic limit (but $\Delta_{\gamma=0}=0$, because $\gamma=0$ corresponds to the unitary case in which the eigenvalues of the Floquet channel have modulus 1).
@@ VII Discussion
First, our results indicate, according to the dissipative diagnostic, that the undoped circuit realizes the strongest intrinsic relaxation in our setting. For any fixed $\gamma>0$, the Liouvillian gap grows linearly with $N$, so relaxation to the steady state accelerates as the system size increases. The mechanism is simple. The unitary dynamics rapidly spreads typical local Pauli operators to extensive weight $w(t)={\cal O}(N)$ on an ${\cal O}(N)$ timescale [Spreading2018]. Once operators are extensive, local noise of strength $\gamma$ damps them at a rate $\sim\gamma w$, i.e. $\sim\gamma N$. In the absence of conserved densities or other slow channels, there is no hydrodynamic mechanism that can produce small decay rates, and the Liouvillian spectrum develops a parametrically large gap.
Second, adding Haar random gates, contrary to our initial intuition based on simulation complexity, weakens intrinsic relaxation, in that the gap no longer grows with system size upon doping. At the level of Pauli dynamics, local randomness strongly mixes the single-site Pauli components, continually regenerating low-weight Pauli string contributions. Under repeated Floquet steps, this regeneration induces return cycles supported entirely within the low-weight sector, each confined to a small contiguous block that translates along the chain, a mechanism absent in the fine-tuned undoped circuit. These low-weight return cycles dominate the slowest nontrivial eigenmodes and set the intrinsic relaxation rate, thereby rendering the gap finite in the thermodynamic limit. Our analytic results precisely show that this crossover from unbounded growth of the gap with system size to a finite Liouvillian gap sets in once the doping density is nonvanishing. Notably, this crossover sets in when the number of doped sites scales linearly with system size, the same regime in which scrambling diagnostics become Haar-like (e.g., OTOCs) [leone2021quantum, magni2025quantum] and higher $t$-moments approach Haar values [haferkamp2022random, harrow2023approximate]. This suggests that relaxation- and scrambling-based diagnostics become nontrivial on parametrically similar doping scales.
@@
Taken together, our finding of a finite-gap onset at linear doping and prior results on Haar-like scrambling suggest parametrically similar scales for relaxation- and scrambling-based diagnostics. Interpreting this comparison, however, requires some care. Most existing results are established in circuit ensembles with a different microscopic architecture, typically at greater depth or in different geometries than the shallow one-dimensional setting analyzed here. Bringing these diagnostics into a single controlled framework therefore calls for extending the present model beyond one dimension and beyond the shallow-depth setting. An important direction for future work is to explore whether the finite-gap scale correlates with approximate unitary $t$-design formation, thereby clarifying how these two diagnostics are related. Ultimately, establishing such links between complementary diagnostics may help develop a more unified picture of quantum chaos beyond diagnostic-specific characterizations.